OpenStudy (anonymous):

HELP...???? Simplify (4xy)(2x2y)(3xy)3. 24x2y3 24x6y5 216x3y3 216x6y5

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@kisstherains @abb0t @ikram002p @ObeySoph▲ @pinker-than-floyd @Broskishelleh

OpenStudy (anonymous):

@jakashaka123 @mathstudent55 @Mathismynemesis @mathmate @mathisblockingmypath

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@jim_thompson5910 @JoannaBlackwelder @jenniferjuice

jimthompson5910 (jim_thompson5910):

What is (3xy)^3 equal to?

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

(3xy)^3 = (3xy) cubed

jimthompson5910 (jim_thompson5910):

(3xy)^3 = (3xy)*(3xy)*(3xy) = ???

OpenStudy (broskishelleh):

D @helpmeplsz